OpenStudy (anonymous):

\[\int\limits_{1}^{\infty}e^{-x} dx\] anyone know how do you evaluate this integal?

OpenStudy (zarkon):

yes

OpenStudy (jhannybean):

\[-e ^{-x}\] evaluated from \[[1,\infty)\]

OpenStudy (anonymous):

Skim through your textbook for 'Improper integral'.

OpenStudy (jhannybean):

So you have \[-e ^{-1}- (-e ^{-\infty}) = e ^{-\infty}-\frac{ 1 }{ e }\]

OpenStudy (anonymous):

oh, thank you. My answer was zero :(.

OpenStudy (primeralph):

Nope

OpenStudy (jhannybean):

What did I do wrong.......

OpenStudy (primeralph):

@klllerbee thinks the answer is 0

OpenStudy (jhannybean):

ohhh.. I see. \[e ^{-\infty}= \frac{ 1 }{ e ^{\infty} }= 0\]

OpenStudy (anonymous):

so the answer is zero?

OpenStudy (primeralph):

nope

OpenStudy (jhannybean):

so now we have \[0-\frac{ 1 }{ e } = -\frac{ 1 }{ e }\]

OpenStudy (anonymous):

so -1/e = 0 ?

OpenStudy (primeralph):

nope, not 0

OpenStudy (jhannybean):

It won't be 0 because you've already integrated and input your limits of integration. the only way \[-\frac{ 1 }{ e } =0\] is if we were still taking the limit. the answer is \[-\frac{ 1 }{ e }\]

OpenStudy (zarkon):

1/e

OpenStudy (anonymous):

so 1/e = -1/e?

OpenStudy (jhannybean):

wai... 0-(-1/e) ..... yes.

OpenStudy (jhannybean):

No i forgot my negative sign. -.- @Zarkon is correct. the answer is (1/e)

OpenStudy (jhannybean):

Thanks @Zarkon

OpenStudy (anonymous):

Oh cool. Zarkon is an android, so he can do it all in his head.

OpenStudy (jhannybean):

That's not true,maybe he's just seen the problem before, or worked it out himself.

OpenStudy (anonymous):

But I can see his picture, he's data

OpenStudy (anonymous):

TY both ;D

OpenStudy (jhannybean):

No problem :)

OpenStudy (zarkon):

you should probably also work on notation too \[\int\limits_{1}^{\infty}e^{-x} dx=\lim_{t\to\infty}\int\limits_{1}^{t}e^{-x} dx\]

OpenStudy (jhannybean):

Awww yiisss.. I hated that substitution.

OpenStudy (jhannybean):

My notation?